a company that sells pocket radios makes a profit of R12000 on the first 500 radios sold. Assuming that all prices remain the same, how much profit do you expect to make on selling 7000 radios?
1
Expert's answer
2020-06-07T13:49:26-0400
"\text {The company makes a profit of 12,000 on the first 500 radios sold.}\\[1 em]

\text {If it sold 7,000 devices at the same price.}\\[1 em]

\text {Now we are looking at how many 500 are there in the 7,000?}\\[1 em]
\frac{7000}{500}=14\\[1 em]
\therefore \text {The expected profit from selling 7,000 radios}\\[1 em]
=12000 \times 14=\textcolor{blue}{168000}\\[2 em] 
\text {Another way}\\\\
\\[2 em]
\begin{aligned}
&12000 \rightarrow &500&\\
&x \rightarrow & 7000&\\
\end{aligned}\\
\therefore x=\frac{12000\times 7000}{500}=\textcolor{blue}{168000}"